2 is added to twice a certain number and the sum is doubled.The result is 10 less than 5 times the original number.find the original number.
Answer by Maths68(1474)

Let x be the number

2(2+2x)=5(x)-10
4+4x=5x-10
4+10=5x-4x
14=x
x=14

Number is 14